how many pounds of water must be evaporated from 50lbs of a 3% salt solution so that the remaining solution will be a 5% salt?
Answer by stanbon(75887)

how many pounds of water must be evaporated from 50lbs of a 3% salt solution so that the remaining solution will be a 5% salt?
----
Equation:
water - water = water
0.97*50 - x = 0.95(50-x)
---
Multiply thru by 100 to get:
97*50 - 100x = 95*50-95x
---
2*50 = 5x
x = 20 lbs (amt. of water to evaporate)
==============
Cheers,
Stan H.